# A quantum random access memory (QRAM) using a polynomial encoding of binary strings

### v1

@@ 3.2 Parallelizing the multiplications : lower T-depth
The set of $N\ell$ Toffolis required to read or write from the memory can be parallelized to have Toffoli-depth 1, with an additional $N\ell$ number of ancillae, as shown in Figure 1(c), where $\ell=1$. During reading the contents of each memory location is copied to an ancilla. Then using CNOTs the parity of these ancillae is stored on the output bus. As explained before, during writing the control and target of the memory locations and output bus are reversed. We can also use the method described in Figure 4 of [DMGM20] in order to parallelize these operations and have Toffoli-depth 1.
Hence, using either of the implementations in [Jon13, Gid18] or [Sel13], accounting for both the computation and uncomputation part, we achieve a T-depth of $2\log_{2}\log_{2}N$, excluding the read/write operation.
@@
Thus, excluding the $N\ell$ memory qubits and $N$ ancillae to parallelize the reading/writing operation, the total number of logical qubits we require is
@@ 3.3 Comparison with previous work
We compare the resource estimates with the parallelized version of bucket-brigade QRAM [DMGM20]. Here we mention that we have compared with the bucket-brigade architecture for the following reasons. First, this has been the most widely studied QRAM and a detailed fault-tolerant resource estimates is available, as in [DMGM20]. Second, its applications are more general than other QRAMs, for example, FF-QRAM [PPR19], EQGAN-QRAM [NZB+22], PQC-based QRAM [PLG23], which have been designed for specific problems and some of them do not scale well. Third, QRAMs as in [PPR19, NZB+22, PLG23] use unitaries like controlled rotations, that are approximately implementable by discrete universal gate sets. The non-Clifford cost like T-count or Toffoli-count varies inversely with the precision of synthesis [GMM22b, Muk24c, Muk24a, Muk24b]. Hence, such designs are more expensive to implement fault-tolerantly. Fourth, many works have used the bucket-brigade QRAM as a basic module and designed more intricate architecture with it, most often to achieve some tradeoffs for particular applications or scenario, for example [XHF+23]. Thus, we can simply replace this module with our $\text{QRAM}_{poly}$ in order to compare the designs.
We compare the Toffoli-count, Toffoli-depth and number of logical qubits. In this way we do not have to worry about the difference in T-count and T-depth due to different implementations of Toffoli. We do not consider resource estimates for the reading/writing operation since this part is same in both the implementations. Without this, the parallel bucket-brigade circuit in [DMGM20] requires $N-2$ compute-uncompute Toffoli pairs, $2N+\log_{2}N$ logical qubits and has Toffoli-depth $\log_{2}N$. From Equations 9, 13 and 15, this implies the following.
@@
In most error-correction schemes, including the most popular surface code, the cost of implementation of the non-Clifford T gate is much more than the cost of the Clifford gates [FMMC12]. In [DMGM20] the following has been taken as a rough estimate for the cost of implementation with the surface code.
@@
In this paper we do not perform a detailed cost analysis for implementation with any quantum error correction code. But from the above equation and Equations 18-20 we can say that our $\text{QRAM}_{poly}$ has much less fault-tolerant cost estimates from previous bucket-brigade architecture. So we expect much better performance in terms of running time.
Remark 3.2.
During parallelizing we use many CNOTs to copy the variables. But these CNOTs are either multi-target ones or/and can be executed in parallel. This is also true for the CNOTs that are used for XOR-ing the monomials. In surface code implementation the cost of implementation of a multi-target CNOT is equal to the cost of a single target CNOT [FMMC12] and has the same execution time. Parallel CNOTs can also be executed in less time. More parallelization of the CNOTs, while adding monomials can be done by using extra ancillae. We remember that the fault-tolerant implementation of CNOT, which is a Clifford, is much less costly than a T or Toffoli gate. So in this paper we optimize the non-Clifford gate count and depth and hence we do not analyse in detail the best CNOT-depth that can be achieved.
@@ 4.1 Quantum look-up-table ($\text{qLUT}_{poly}$)
In the CSWAP architecture for qLUT [LKS24] the authors combine a QROM and a specific QRAM. The QROM is implemented with a set of multi-controlled-NOT gates. The QRAM is implemented with a number of Fredkin or controlled-SWAP unitaries. Each controlled-SWAP can be implemented with a Toffoli and CNOT. The contents of the selected memory location is always obtained on some specified qubits. This qLUT has a T-count of $O(\sqrt{N})$, T-depth $O(\sqrt{N})$ and number of qubits $O(\sqrt{N})$.
@@
Our CNOT cost is primarily dominated by the step where we copy a subset of memory locations selected by the first QRAM. We require $\ell N_{1}N_{2}=\ell N$ CNOTs at this stage. But again this is a group of $N_{1}$ multi-target CNOTs, where each has $N_{2}$ target. So in surface code implementation the execution time is equivalent to the time of execution of $N_{1}$ logical CNOTs. This cost is the same as that in [LKS24].
@@ 4.2 Toffoli-count optimization of quantum circuits
The polynomial encoding can be used to optimize the number of Toffolis required to implement groups of multi-controlled-X gates or mixed polarity multiple control Toffolis (MPMCTs) [DMGM20]. For example, we want to implement a circuit that flips a qubit to |$\ket{1}$ for a subset, $\mathcal{S}\subseteq\{0,1\}^{n}$, of $n$-bit strings. These types of circuits also represent a kind of QROM. These can be used to select a subset of addresses and implement certain operations on those locations. We optimize the Toffoli-count of such circuits using the following procedure, which we call TOFFOLI-OPT-POLY.
@@
Here we observe that a sequence of multi-controlled-X gates represents a Boolean function which is a sum of product terms. We can find its ESOP (Exclusive Sum-of-Products) expression using tools like EXORCISM [SP93, SP96, MP01]. Then factoring this expression we can implement the Boolean expression. We can also use the algorithm in [SSP13] which first computes the ESOP, and then breaks the expression into common cofactors, which are reversibly synthesized. For example, in [DMGM20] the authors mentioned that the circuit in Figure 4(a) can be implemented with 2 Toffolis. This is more than the Toffoli-count we get. We can also implement each multi-controlled-NOT-X gate using the decomposition given in [HLZ+17]. Each $C^{n}X$ i.e. an $n$-qubit-controlled-X gate can be implemented with $n-1$ Toffolis and an additional $n-1$ ancillae. But this will give more Toffoli-count than our implementation in Figure 4(b).
@@ 5 Discussions and Conclusion
In this paper we develop a new design for quantum random access memory (QRAM), using a polynomial encoding of the bit strings specifying the address of memory locations. Each memory location is associated to a unique polynomial, such that with a specific bit string only one of these polynomials has value 1 mod 2, and the rest have value 0 mod 2. We prove a number of properties of these encoding polynomials, with which we implement a Clifford+T circuit for our $\text{QRAM}_{poly}$. Each monomial in these polynomials can be implemented with Toffolis, the only non-Clifford gate we use. We show that $\text{QRAM}_{poly}$ has T-count $O(N-\log N-1)$, T-depth $O(\log\log N)$ and uses $O(N)$ logical qubits. We observe that with our design of $\text{QRAM}_{poly}$ we achieve (Table 1) an exponential improvement in T-depth, while reducing T-count and keeping the number of logical qubits requirement the same with respect to the previous state-of-the-art bucket brigade architecture [GLM08a, DMGM20]. The product of T-depth and number of logical qubits is taken as a rough estimate for fault-tolerant implementation with the surface code. Hence an exponential improvement in T-depth while keeping the qubit-count same and in fact reducing the T-count, is expected to yield a better performance in the running time [Fow12, FMMC12].
Using two such $\text{QRAM}_{poly}$ we implement a quantum look-up table ($\text{qLUT}_{poly}$) that has T-count $O(\sqrt{N})$, T-depth $O(\log\log N)$ and uses $O(\sqrt{N})$ logical qubits. With our quantum look-up-table circuit $\text{qLUT}_{poly}$ we achieve (Table 1) a double exponential improvement in T-depth over the previous state-of-the-art CSWAP architecture for qLUT [LKS24], while the T-count and qubit-count are asymptotically same.
@@
Using the polynomial encoding, we develop a method (TOFFOLI-OPT-POLY) to optimize the Toffoli-count of quantum circuits, especially those using multi-controlled-NOT gates. Since such circuits represent a sum-of-product (SOP) form of Boolean function, so these encodings can also have potential application in optimizing Boolean ESOP expressions, similar to the algorithms in [SP93, SP96, MP01, SSP13].
This paper mainly aims at optimizing the non-Clifford gate complexity like T-count and T-depth, which is important from a fault-tolerant perspective. It will be interesting to study the noise resilience of the circuits developed in this paper and compare it with existing architectures, as done in [AGJO+15, HLGJ21]. It will also be interesting to study the mapping overhead of these circuits in different architectural layouts like 2D grid, as done in [XHF+23, ZSL24]. Once we know the trade-offs, we can aim at developing different hybrid designs with these new and existing circuits, so that we can take advantage of the various designs in different scenarios. Another direction can be to investigate more applications for the polynomial encoding developed in this paper, for example, in Boolean ESOP expression optimization (as mentioned earlier), design of efficient application-specific QRAMs, as in [PPR19, NZB+22, PLG23]. We leave these questions for future research.
@@ Acknowledgement
The author thanks Nathan Wiebe for helpful discussions.

### v2

@@ 3.3 Parallelizing the multiplications : lower T-depth
The set of $N\ell$ Toffolis required to read or write from the memory can be parallelized to have Toffoli-depth 1. A simple method using an additional $N\ell$ number of ancillae, has been shown in Figure 2(c), where $\ell=1$. During reading, the contents of each memory location is copied to an ancilla. Then using CNOTs the parity of these ancillae is stored on the output bus. As explained before, during writing the control and target of the memory locations and output bus are reversed. In order to parallelize these operations and have Toffoli-depth 1, we can also use the method described in Figure 4 of [29] or in [71], the latter has exponentially less qubit-count.
Hence, using either of the implementations in [68, 69] or [70], accounting for both the computation and uncomputation part, we achieve a T-depth of $2\log_{2}\log_{2}N$, excluding the read/write operation.
@@
Thus, excluding the memory qubits and ancillae required to parallelize the reading/writing operation, the total number of logical qubits we require is
@@ 3.4 Comparison with previous work
We compare the resource estimates with the parallelized version of bucket-brigade QRAM [29]. Here we mention that we have compared with the bucket-brigade architecture for the following reasons. First, this has been the most widely studied QRAM and a detailed fault-tolerant resource estimates is available, as in [29]. Second, its applications are more general than other QRAMs, for example, FF-QRAM [40], EQGAN-QRAM [72], PQC-based QRAM [49], which have been designed for specific problems and some of them do not scale well. Third, QRAMs as in [40, 72, 49, 73, 74] use unitaries like controlled rotations, that are approximately implementable by discrete universal gate sets. The non-Clifford cost like T-count or Toffoli-count varies inversely with the precision of synthesis [75, 62, 61, 63]. Hence, such designs are more expensive to implement fault-tolerantly. Fourth, many works have used the bucket-brigade QRAM as a basic module and designed more intricate architecture with it, most often to achieve some tradeoffs for particular applications or scenario, for example [36]. Thus, we can simply replace this module with our $\text{QRAM}_{poly}$ in order to compare the designs.
We compare the Toffoli-count, Toffoli-depth and number of logical qubits. In this way we do not have to worry about the difference in T-count and T-depth due to different implementations of Toffoli. We do not consider resource estimates for the reading/writing operation since this part can be implemented and optimized in a similar fashion and hence its cost can be regarded the same. Without this, the parallel bucket-brigade circuit in [29] requires $N-2$ compute-uncompute Toffoli pairs, $2N+\log_{2}N$ logical qubits and has Toffoli-depth $\log_{2}N$. From Equations 9, 13 and 15, this implies the following.
@@
In most error-correction schemes, including the most popular surface code, the cost of implementation of the non-Clifford T gate is much more than the cost of the Clifford gates [64]. In [29] the following has been taken as a rough estimate for the cost of implementation with the surface code.
@@
Thus from Equations 17-19 we can say that our $\text{QRAM}_{poly}$ has much less fault-tolerant cost estimates from previous bucket-brigade architecture. So we expect much better performance in terms of running time.
For illustration, we consider an implementation of our $\text{QRAM}_{poly}$ with the surface code, and estimate resource requirements following the procedures described in [64] and [37]. Consider the number of qubits $n=36$, which corresponds to 8 GB of classical data. Using logical AND gadgets [69], Toffoli can be implemented with 4 T gates with a T-depth 2 and if there is a compute-uncompute Toffoli pair we do not require any T gate for the uncomputation part. Hence, from Equation 9, assuming $\ell=1$, the T-count is
$\mathcal{T}\leq 4(2N-\log n-1)=5.4975\times 10^{11}$ (21)
and the T-depth is
$\mathcal{T}_{d}\leq 2(\lceil\log n\rceil+1)=14.$ (22)
Each T gate requires one magic state (|$\ket{A_{L}}$). For the above T-count the output error rate for state distillation should be no greater than
$p_{out}=\frac{1}{\mathcal{T}}\approx 1.81899^{-12}.$ (23)
Now we use Algorithm 4 in [37] in order to calculate the distance of the surface code. Assuming a magic state injection error rate $p_{in}=10^{-4}$, a per-gate error rate $p_{g}=10^{-5}$, the stated algorithm suggests 2 layers of distillation with distances $d_{1}=10$, $d_{2}=5$. The first stage of distillation consumes $16$ logical qubits and the second stage consumes $16\times 15=240$ logical qubits in order to generate a single magic state.
The input states of the logical qubits in the second layer are encoded on a distance $d_{2}=5$ code that uses $N_{p2}=2.5\times 1.25\times d_{2}^{2}\approx 79$ physical qubits per logical qubit. The total footprint of the distillation circuit is then $18750240\times N_{p2}=$ 18750 physical qubits. This round of distillation is completed in $\sigma_{2}=10d_{2}=50$ surface code cycles.
The first or top layer requires a $d_{1}=10$ surface code, for which a logical qubit takes $N_{p1}=2.5\times 1.25\times d_{1}^{2}\approx 313$ physical qubits. So the total number of physical qubits required is $500016\times N_{p1}=$ 5000, with the round of distillation completed in $\sigma_{1}=10d_{1}=100$ surface code cycles.
The concatenated distillation scheme is performed in $\sigma=\sigma_{1}+\sigma_{2}=150$ surface code cycles. Since the first or top layer has lower footprint than the second or bottom layer, distillation can potentially be pipelined to produce
$\frac{150\times 18750}{100\times 5000+50\times 18750}\approx 2$
magic states in parallel. The physical qubits in the second layer is reused. Let $t_{sc}=200ns$ is a surface code cycle time. Then 2 magic states can be produced every $\sigma\times t_{sc}=30\times 10^{-6}$ s. We require $\frac{\mathcal{T}}{\mathcal{T}_{d}}=\frac{5.4975\times 10^{11}}{14}\approx 3.9\times 10^{10}$ magic states per layer or depth of T-gates. We produce these many magic states in parallel for each layer. Due to parallelization the number of physical qubits required is
$\frac{1}{2}\times 3.9\times 10^{10}\times 18750\approx 3.66\times 10^{14}$ (24)
and the time taken is
$14\times 30\times 10^{-6}s=4.2\times 10^{-4}s.$ (25)
In surface code implementation the cost of implementation of a multi-target CNOT is equal to the cost of a single target CNOT [64] and has the same execution time. So we consider a multi-target CNOT as one logical CNOT. Each Toffoli can be implemented with 2 CNOTs, 2 multi-target CNOTs, 1 H and 1 S gate [69]. Thus we can upper bound the number of logical Cliffords by $7N=7\times 2^{36}=4.81\times 10^{11}$. The overall error rate of the Cliffords should therefore be less than $\frac{1}{7N}\approx 2.08\times 10^{-12}$. To compute the required distance, we seek the smallest $d$ that satisfies the inequality
$\left(\frac{p_{in}}{0.0125}\right)^{\frac{d+1}{2}}<2.08\times 10^{-12}$
and find this to be $d=11$. The number of physical qubits required to encode the Cliffords is at most $2\times 2^{36}\times 2.5\times 1.25\times 11^{2}\approx 5.197\times 10^{13}$. Overall, we require approximately $3.66\times 10^{14}+5.197\times 10^{13}\approx 4.18\times 10^{14}$ physical qubits to encode the complete $\text{QRAM}_{poly}$.
Roughly, we expect to perform $\frac{7N}{2N\times\mathcal{T}_{d}}=\frac{7}{2\times 14}\approx 0.25$ logical Clifford operations per qubit per layer of T-depth. Since each logical CNOT takes 2 surface code cycles and we require $\sigma=150$ surface code cycles per T-depth, so the overall time is dominated by the time for implementing the non-Clifford T-gates. Thus, the time for implementing one memory read/write operation with $\text{QRAM}_{poly}$ is approximately $4.2\times 10^{-4}$s and utilizes about $4.18\times 10^{14}$ physical qubits.
The QRAM in [29] requires $1.5\times 10^{15}$ physical qubits and is implemented in approximately $2.13\times 10^{-3}$s, using the same surface code parameters and doing a similar analysis. We remark that design considerations can vary. We can further reduce the time by parallelizing the magic state factories even more. This will increase the number of physical qubits.
@@ 4.1 Quantum look-up-table ($\text{qLUT}_{poly}$)
In the CSWAP architecture for qLUT [58] the authors combine a QROM and a specific QRAM. The QROM is implemented with a set of multi-controlled-NOT gates. The QRAM is implemented with a number of Fredkin or controlled-SWAP unitaries. Each controlled-SWAP can be implemented with a Toffoli and CNOT. The contents of the selected memory location is always obtained on some specified qubits. This qLUT has a T-count of $O(\sqrt{N})$, T-depth $O(\sqrt{N})$ and number of qubits $O(\sqrt{N})$.
@@
Our CNOT cost is primarily dominated by the step where we copy a subset of memory locations selected by the first QRAM. We require $\ell N_{1}N_{2}=\ell N$ CNOTs at this stage. But again this is a group of $N_{1}$ multi-target CNOTs, where each has $N_{2}$ target. So in surface code implementation the execution time is equivalent to the time of execution of $N_{1}$ logical CNOTs. This cost is the same as that in [58].
@@ 4.2 Toffoli-count optimization of quantum circuits
The polynomial encoding can be used to optimize the number of Toffolis required to implement groups of multi-controlled-X gates or mixed polarity multiple control Toffolis (MPMCTs) [29]. For example, we want to implement a circuit that flips a qubit to |$\ket{1}$ for a subset, $\mathcal{S}\subseteq\{0,1\}^{n}$, of $n$-bit strings. These types of circuits also represent a kind of QROM. These can be used to select a subset of addresses and implement certain operations on those locations. We optimize the Toffoli-count of such circuits using the following procedure, which we call TOFFOLI-OPT-POLY.
@@
Here we observe that a sequence of multi-controlled-X gates represents a Boolean function which is a sum of product terms. We can find its ESOP (Exclusive Sum-of-Products) expression using tools like EXORCISM [76, 77, 78]. Then factoring this expression we can implement the Boolean expression. We can also use the algorithm in [79] which first computes the ESOP, and then breaks the expression into common cofactors, which are reversibly synthesized. For example, in [29] the authors mentioned that the circuit in Figure 5(a) can be implemented with 2 Toffolis. This is more than the Toffoli-count we get. We can also implement each multi-controlled-NOT-X gate using the decomposition given in [80]. Each $C^{n}X$ i.e. an $n$-qubit-controlled-X gate can be implemented with $n-1$ Toffolis and an additional $n-1$ ancillae. But this will give more Toffoli-count than our implementation in Figure 5(b).
@@ 5 Discussions and Conclusion
In this paper we develop a new design for quantum random access memory, using a polynomial encoding of the bit strings specifying the address of memory locations. We implement a Clifford+T circuit for our $\text{QRAM}_{poly}$ and show that $\text{QRAM}_{poly}$ has T-count $O(N-\log N-1)$, T-depth $O(\log\log N)$ and uses $O(N)$ logical qubits. Thus with our design of $\text{QRAM}_{poly}$ we achieve an exponential improvement in T-depth, while reducing T-count and keeping the number of logical qubits requirement the same with respect to the previous state-of-the-art bucket brigade architecture [20, 29]. We illustrate that when encoded with the surface code [65, 64], in order to perform one memory read/write operation, $\text{QRAM}_{poly}$ takes less time and uses much less number of physical qubits. Using two such $\text{QRAM}_{poly}$ we implement a quantum look-up table ($\text{qLUT}_{poly}$) that has T-count $O(\sqrt{N})$, T-depth $O(\log\log N)$ and uses $O(\sqrt{N})$ logical qubits. With our quantum look-up-table circuit $\text{qLUT}_{poly}$ we achieve (Table 1) a double exponential improvement in T-depth over the previous state-of-the-art CSWAP architecture for qLUT [58], while the T-count and qubit-count are asymptotically same.
In our designs reduction in non-Clifford gate count comes at the cost of an increase in the CNOT gate count. The latter is a Clifford gate and in most error correction schemes the cost of implementing a Clifford is much less than the cost of implementing a non-Clifford. Thus, in our illustration we obtained better performance compared to previous QRAM designs. But CNOT, being a multi-qubit gate is more error prone than single-qubit gates like T. Even for connectivity constrained architectures (especially of the NISQ era) implementing a multi-qubit gate becomes more costly because it often needs a number of intermediate CNOT or SWAP gates, thus increasing the total gate count [81].
The problem of studying the noise-resilience of QRAM is an active research problem [27, 28], especially for the pre-fault-tolerant regime. And we believe it is beyond the scope of this current work because this paper exclusively focuses on performance improvements in the fault-tolerant regime. Often metrics and design considerations in these two regimes differ and hence they are studied separately. A detail analysis of the noise-resilience of $\text{QRAM}_{poly}$ is left for future work. It will also be interesting to study the mapping overhead of these circuits in different architectural layouts like 2D grid, as done in [36, 59]. We expect to find trade-offs between the CNOT count and error rate or sparsity of the underlying graph. We can also aim at developing different hybrid designs with these new and existing circuits, so that we can take advantage of the various designs in different scenarios.
@@
Using the polynomial encoding, we develop a method (TOFFOLI-OPT-POLY) to optimize the Toffoli-count of quantum circuits, especially those using multi-controlled-NOT gates. Since such circuits represent a sum-of-product (SOP) form of Boolean function, so these encodings can also have potential application in optimizing Boolean ESOP expressions, similar to the algorithms in [76, 77, 78, 79]. Some of these classical algorithms have inspired methods for reversible quantum logic synthesis [73], which in turn have been an integral part of the design of quantum oracles for important algorithms like Grover’s search. Thus these encoding polynomials may be used for reversible quantum logic synthesis. In the future we aim to investigate this avenue and the application of the polynomial encodings towards the design of algorithm-specific oracles and application-specific QRAMs, as in [40, 72, 49, 73, 74].
@@ Acknowledgement
The author thanks Nathan Wiebe for helpful discussions. The author also thanks the anonymous reviewers whose helpful comments have helped us improve our manuscript significantly. The author acknowledges funding from the NSERC discovery program.
Author contributions
The ideas, implementations and preparation of the manuscript was done by P.Mukhopadhyay.
Data availability
All relevant data are included in this manuscript.
Competing interests
There are no competing interests to declare.